# Motion Planning Diffusion: Learning and Adapting Robot Motion Planning with Diffusion Models

### v1

@@ 6             ${\bm{\mu}}_{i}^{0}={\bm{\mu}}_{i}$
                   $\triangleright$ compute costs gradient eqs. 28 and 29
@@ IV-A Environments, Tasks and Datasets
We consider a set of tasks that are representative of the challenges faced in robot motion planning, from simple to complex, including a $2$D point mass navigating an environment with simple obstacles, a $2$D point mass navigating a narrow passage, a $2$-link planar robot, a $4$-link planar robot, and a $7$-dof robot arm manipulator in two settings: one illustrative environment with collision spheres; and a warehouse environment with a table and two shelves. The tasks are displayed in fig. 5, and the naming convention is [Environment]-[Robot]. In all tasks, the goal is to move the robot from an initial to a goal configuration – either a desired joint position or an end-effector pose – while avoiding collisions with the environment.
@@
Figure 5: The environments used for the motion planning experiments, with different robot models and increasing complexity. In $2$D tasks the robots are depicted in blue (as a dot for RobotPointMass2D). The bottom row shows the tasks with additional obstacles (in red), which are not present during training.
For each task, we generate a dataset of collision-free paths with the OMPL library Python bindings [105], along with PyBullet [106] to setup the environment and perform collision checking. We randomly sample multiple contexts and use RRT Connect [5] to generate paths, which are afterward shortcutted and smoothed. Moreover, we generate an additional path by reverse-connecting the goal to the start configuration. Alternatively, if we know the task requires goal poses to be located in some regions of the workspace, we can specify the goal pose and use inverse kinematics to obtain the goal joint configuration. For example, in the EnvWarehouse-RobotPanda task, the goal pose is located on top of the table or on the shelves, and we sample paths from any initial configuration to those goal poses. There are two reasons for using RRT Connect to generate data. First, because it is faster than optimal planners like RRT, and since we optimize these paths further using defined cost functions, we trade optimality for speed. Second, it is probabilistically complete, and its vertices cover all the configuration-free space (Corollary $1$ of [5]), thus finding multiple modes for reaching the goal from the start configuration. The number of paths, which equals the number of contexts, generated per environment can be found in table II. Since we are planning in joint space, the number of points needed to cover the whole space increases exponentially with the number of joints. Thus, a sufficient amount of data is needed to counter the curse of dimensionality. E.g., for the discretization of $100$ points per joint, a $4$-dof robot arm requires $10100^{4}=$ 10M points to cover the discretized space. When training the diffusion model prior, we sample a context and path from the dataset and fit a B-spline to the path using the splprep function from the SciPy library [107]. The number of B-spline control points is a hyperparameter and is chosen so the path almost does not collide with the environment. Note that we do not need to be 100% collision-free, as collision-avoidance costs are used during optimization.
@@ IV-B Baselines
  • To assess the benefits of using a diffusion model, we use a baseline that learns a Conditional Variational Autoencoder (CVAE) [109] to model the trajectory distribution. It shares the same U-Net architecture as the denoising function ${\bm{\epsilon}}_{{\bm{\theta}}}$, but at the lowest level of the U-Net the embedding is projected into a $32$-dimensional latent space that encodes the mean and standard deviation of a Gaussian distribution. The optimized loss function is a sum of L2222-norm of the predicted and ground-truth trajectory and the KL-divergence between the latent space posterior and a standard Gaussian distribution weighted by $\beta=0.1$.
@@ IV-D General Results in Simulation
Figure 6: Performance metrics for different algorithms on the tasks from fig. 5. The results report the mean value of sampling $100$ contexts from the test set and optimizing $100$ trajectories per context. The columns under training environment show the results without additional objects. In $2D$ environments, the errors in end-effector position and orientation are always 00 since a joint goal is provided.
@@
In this experiment, we report the results using all motion planning costs. Here, we want to evaluate how the methods perform in terms of collisions and reaching the desired end-effector goal poses. Therefore, we use a large time duration of $10$ seconds to prevent optimization over velocities and accelerations. We sample $100$ contexts from the test set and optimize $100$ trajectories per context in batch. The results are displayed in fig. 6.
@@
To answer Q1, we observe the results under training environment since they show how well the generative priors model the data distribution. They show that the diffusion-based models (Dprior, Dprior+Cost, MPD) obtain success rates that match or surpass the baselines while keeping a high diversity in generated trajectories. When comparing CVAE and Dprior, their success rate and validity fraction are close, but the diversity score is higher for the diffusion-based models, in particular for higher-dimensional environments (using RobotPanda). For instance, in the EnvWarehouse-RobotPanda, both CVAE and Dprior achieve $97\%$ success rate, but Dprior shows more variability ($50.9$ vs. $60.3$). Moreover, Dprior achieves smaller errors in the desired end-effector goal position and orientation ($7.4$cm vs. $5.2$cm, $10.5^{\circ}$ vs. $6.7^{\circ}$), which due to model approximations are not 00. These errors are improved during cost optimization. E.g., MPD achieves in this task position and orientation errors of $2.00$cm and $1.1^{\circ}$, respectively.
@@
To answer Q2, we look at the results when using newer obstacles in the scene under additional objects columns. The success rate and validity fraction drop for all methods, in comparison to the training environment, which is expected, in particular for the prior methods, CVAE and Dprior, which do not have knowledge of the new obstacles. For instance, in the EnvNarrowPassageDense2D-RobotPointMass2D task, Dprior’s success rate dropped from $95\%$ to $8\%$. This is a task where an informed prior is particularly useful since traversing from the top to the bottom of the environment needs to be done through a very small passage. We focus now on the higher-dimensional task EnvWarehouse-RobotPanda, as the conclusions transfer to other tasks. In the presence of new obstacles, all methods that perform cost optimization improve the success rate and validity fraction in comparison to using only the prior. For example, MPD increases the validity fraction of Dprior from $18.4$ to $73.5\%$, which is the largest among all methods for this task. Moreover, MPD also shows more diversity in the valid trajectories in comparison to baselines that first sample from the prior and then optimize the cost ($66.0$ for Dprior+Cost vs. $74.8$ for MPD). Note that these results are more pronounced for the additional objects tasks, and a discussion for why this happens is found in section IV-E. When comparing MPD with an uninformed prior GPprior+Cost, we observe a success rate increase from $67.8\%$ to $96.0\%$, which supports the claim that the diffusion prior is useful to guide the trajectories through collision-free regions. In terms of diversity, we expect the GPprior to provide the largest value since this is an uninformed prior. MPD’s diversity is the closest to the GPprior one or slightly larger for the EnvWarehouse-RobotPanda task. These results show that MPD produces a higher percentage of collision-free trajectories (and inside joint limits) while being diverse. It is also important to note the effect of optimizing the task cost. Methods that optimize this cost produce trajectories resulting in end-effector pose errors much smaller than the ones obtained just sampling from the priors. E.g., in the EnvWarehouse-RobotPanda task, the end-effector error in orientation drops from $7.1$ to $1.2$ between Dprior and MPD.
@@
[…]
Figure 10: The figures illustrate the benefits of sampling with MPD instead of first sampling from the prior and then optimizing the cost function. LABEL:sub@fig:EnvEmpty2D_demonstrations The robot is a $2$D point mass in an environment without any obstacles, and demonstration trajectories are obtained by sampling from a Gaussian Process prior. At inference, an obstacle (in red) is inserted into the environment. LABEL:sub@fig:EnvEmpty2D_diffusion_prior Samples from the learned diffusion prior. LABEL:sub@fig:EnvEmpty2D_diffusion_prior_then_guide Results of first sampling from the prior and then optimizing the collision cost function. LABEL:sub@fig:EnvEmpty2D_mpd Samples from the posterior distribution generated with MPD. The second row illustrates how MPD’s denoising process moves from Gaussian noise to clean trajectories while avoiding the red obstacle. Trajectories that are in collision with the red obstacles are shown in black, and collision-free trajectories in orange.
In this section, we answer Q3. One important aspect when using cost-guided diffusion in the context of motion planning is to understand in what situations can sampling from the denoising posterior distribution eq. 8 be beneficial, compared to first sampling from the diffusion prior and then optimizing the cost function starting from the proposed samples. In fig. 10 we show a simple example to motivate this difference. We consider a $2$D point mass in an environment without any obstacle (fig. 10a), generate a small set of demonstrations with a Gaussian Process trajectory, and learn a trajectory diffusion prior. At inference we add a new obstacle, the red square, and do optimization using the object collision cost. The figures show two aspects of MPD. First, while sampling from the diffusion prior and then optimizing the cost fails to provide collision-free trajectories (fig. 10c), MPD can obtain a larger percentage of those (fig. 10d). This phenomenon happens because once a large portion of the trajectory is in collision with the obstacle, it becomes more difficult to remove it from collision. Second, in fig. 10d, we notice how trajectory samples from MPD evolve during the denoising process. The evolution shows that sampling from the posterior distribution moves the samples towards regions of the configuration space where the posterior has higher probability density, thus generating samples that are collision-free but close to the prior distribution. Therefore, the benefits of sampling from the posterior are more diverse and collision-free trajectories, which can be observed in fig. 6 by comparing Dprior+Cost and MPD results, especially in the environments with additional objects.
@@ IV-F Impact of the B-spline Parametrization
One of this work’s contributions is performing diffusion in the space of B-spline coefficients instead of waypoints. Hence, to answer Q4, we compare both representations. The benefits have been detailed in section III-E, but perhaps the main one is generating smooth trajectories. As the velocity and acceleration of a linear waypoint trajectory representation are computed with central finite differences, they are constant between waypoints and lead to large jumps. There are two ways to get smoother trajectories with this representation. First, we can increase the number of waypoints, as commonly done in optimization-based algorithms. Second, after running the denoising process, we can do extra optimization steps to smoothen and optimize the trajectory to remain collision-free and within joint limits while minimizing the task cost. Both of these options have drawbacks in terms of additional computations, e.g., in the first case, the denoising network needs to process larger inputs. We performed an ablation experiment where we replaced the B-spline with waypoints. We used the same number of waypoints as B-spline coefficients, which guarantees that the denoising network size and computations are similar. As expected, there are no major differences across all metrics except for smoothness values  [^1] Smoothness is measured as the sum of accelerations along the trajectory., which are better (lower) for the B-spline compared to the sparse waypoint representation. To better grasp the difference visually, fig. 11 shows the distribution of smoothness values obtained with MPD across all contexts and tasks. The waypoint representation consistently leads to worse smoothness. Figure 12 shows a visual comparison in the EnvSimple2D-RobotPointMass2D tasks. The figures show that even though the waypoint representation can generate valid trajectories, they are not smooth, as can be seen in the straighter path in the environment plot and in the velocity and acceleration time profiles. These types of trajectories are harder for a robot controller to follow and can lead to jerky movements.
@@ IV-G Learning and Adapting from Human Demonstrations
Learning the prior distribution on trajectories is agnostic to the expert demonstrations. Previously, we assumed that trajectories would be generated by a sampling-based path planner. However, in real-world tasks, it is sometimes desirable and easier for a human to provide demonstrations via kinesthetic teaching of what movement she would like the robot to execute. Therefore, to answer Q5, in this experiment, we learn the trajectory distribution directly from human demonstrations, showing that MPD can be used in this scenario as well. Using the EnvWarehouse-RobotPanda environment, we consider the task of moving one item (a tea box) from several locations in the small to the big shelf and letting a user perform multiple demonstrations from random start and goal poses by moving the robot in gravity compensation with kinesthetic teaching from regions on the small shelf to the big shelf, generating approximately $100$ trajectories. These are augmented using small noise values to perform a total of approximately $4$k trajectories. Examples of human-demonstrated trajectories are shown in fig. 13. The diffusion prior model is then trained on joint space trajectories similarly and with the same hyperparameters as the simulation experiments.

### v3

@@ 6    ${\bm{\mu}}_{i}^{0}={\bm{\mu}}_{i}$
         $\triangleright$ compute costs gradient eqs. 27 and 28
@@ IV-A Environments, Tasks and Datasets
We consider a set of tasks that are representative of the challenges faced in robot motion planning, from simple to complex, including a 22D point mass navigating an environment with simple obstacles, a 22D point mass navigating a narrow passage, a $2$-link planar robot, a $4$-link planar robot, and a $7$-dof robot arm manipulator in two settings: one illustrative environment with collision spheres; and a warehouse environment with a table and two shelves. The latter aims to motivate the applicability of our method in a real-world task, such as shelf rearrangements in distribution facilities, where shelves are static, but at deployment new objects can be placed on the table or shelves. The tasks are displayed in fig. 5, with naming convention [Environment]-[Robot]. The goal is to move the robot (without collisions) from start to goal configurations – a desired joint position or an end-effector pose.
@@
Figure 5: The environments used for the motion planning experiments, with different robot models and increasing complexity. In 22D tasks the robots are depicted in blue (as a dot for RobotPointMass2D). The bottom row shows the tasks with additional obstacles (in red), which are not present during training.
For each task, we generate a dataset of collision-free paths with the OMPL library Python bindings [105], along with PyBullet [106] to setup the environment and perform collision checking. We randomly sample multiple contexts and use RRT Connect [5] to generate paths, which are afterward shortcutted and smoothed. Moreover, we generate an additional path by reverse-connecting the goal to the start configuration. Alternatively, if we know the task requires goal poses to be located in some regions of the workspace, we can specify the goal pose and use inverse kinematics to obtain the goal joint configuration. For example, in the EnvWarehouse-RobotPanda task, the goal pose is located on top of the table or on the shelves, and we sample paths from any initial configuration to those goal poses. There are two reasons for using RRT Connect to generate data. First, because it is faster than optimal planners like RRT, and since we optimize these paths further using defined cost functions, we trade optimality for speed. Second, it is probabilistically complete, and its vertices cover all the configuration-free space (Corollary $1$ of [5]), thus finding multiple modes for reaching the goal from the start configuration. The number of paths, which equals the number of contexts, generated per environment can be found in table II. Since we are planning in joint space, the number of points needed to cover the whole space increases exponentially with the number of joints. Thus, a sufficient amount of data is needed to counter the curse of dimensionality. E.g., for the discretization of $100$ points per joint, a $4$-dof robot arm requires $100^{4}=100$M points to cover the discretized space. When training the diffusion model prior, we sample a context and path from the dataset and fit a B-spline to the path using the splprep function from the SciPy library [107]. The number of B-spline control points is a hyperparameter and is chosen so the path almost does not collide with the environment. Note that we do not need to be 100% collision-free, as collision-avoidance costs are used during optimization.
@@ IV-B Baselines
  • To assess the benefits of using a diffusion model, we use a baseline that learns a Conditional Variational Autoencoder (CVAE) [109] to model the trajectory distribution. It shares the same U-Net architecture as the denoising function ${\bm{\epsilon}}_{{\bm{\theta}}}$, but at the lowest level of the U-Net the embedding is projected into a $32$-dimensional latent space that encodes the mean and standard deviation of a Gaussian distribution. The optimized loss function is a sum of L22-norm of the predicted and ground-truth trajectory and the KL-divergence between the latent space posterior and a standard Gaussian distribution weighted by $\beta=0.1$.
@@ IV-D General Results in Simulation
Figure 6: Performance metrics for different algorithms on the tasks from fig. 5. The results report the swarm plot of sampling $100$ contexts from the test set and optimizing $100$ trajectories per context. The cross represents the mean value. The columns under training environment show the results without additional objects. In $2D$ environments, the errors in end-effector position and orientation are always 0 since a joint goal is provided.
@@
In this experiment, we report the results using all motion planning costs. Here, we want to evaluate how the methods perform in terms of collisions and reaching the desired end-effector goal poses. Therefore, we use a large time duration of $10$ seconds to prevent optimization over velocities and accelerations. We sample $100$ contexts and optimize $100$ trajectories per context. The results are displayed in fig. 6.
@@
To answer Q1, we observe the results under training environment since they show how well the generative priors model the data distribution. They show that the diffusion-based models (Dprior, Dprior+Cost, MPD) obtain success rates that match or surpass the baselines while keeping a high diversity in generated trajectories. When comparing CVAE and Dprior, their success rate and validity fraction are close, but the diversity score is higher for the diffusion-based models, in particular for higher-dimensional environments (using RobotPanda). For instance, in the EnvWarehouse-RobotPanda, both CVAE and Dprior achieve $97\%$ mean success rate, but Dprior shows more mean variability ($50.9$ vs. $60.3$). Moreover, Dprior achieves smaller mean errors in the desired end-effector goal position and orientation (7.47.4cm vs. 5.25.2cm, $10.5^{\circ}$ vs. $6.7^{\circ}$), which due to model approximations are not 0. These errors are improved during cost optimization. Both diffusion-based models, Dprior+Cost and MPD, achieve similar mean position and orientation errors, e.g., in this task, 2.002.00cm and $1.1^{\circ}$.
@@
To answer Q2, we look at the results when using newer obstacles in the scene under additional objects columns. The success rate and validity fraction drop for all methods, in comparison to the training environment, which is expected, in particular for the prior methods, CVAE and Dprior, which do not have knowledge of the new obstacles. For instance, in the EnvNarrowPassageDense2D-RobotPointMass2D task, Dprior’s mean success rate dropped from $95\%$ to $8\%$. This is a task where an informed prior is particularly useful since traversing from the top to the bottom of the environment needs to be done through a very small passage. We focus now on the higher-dimensional task EnvWarehouse-RobotPanda. In the presence of new obstacles, all methods that perform cost optimization improve the success rate and validity fraction in comparison to using only the prior. For example, MPD increases the mean validity fraction of Dprior from $18.4$ to $73.5\%$, which is the largest among all methods for this task. Moreover, MPD also shows more mean diversity in the valid trajectories in comparison to baselines that first sample from the prior and then optimize the cost ($66.0$ for Dprior+Cost vs. $74.8$ for MPD). Note that these results are more pronounced for the additional objects tasks, and a discussion for why this happens is found in section IV-E. When comparing MPD with an uninformed prior GPprior+Cost, we observe a mean success rate increase from $67.8\%$ to $96.0\%$, which supports the claim that the diffusion prior is useful to guide the trajectories through collision-free regions. In terms of diversity, we expect the GPprior to provide the largest value since this is an uninformed prior. MPD’s diversity is the closest to the GPprior one or slightly larger for the EnvWarehouse-RobotPanda task. These results show that MPD produces a higher percentage of collision-free trajectories (and inside joint limits) while being diverse. It is also important to note the effect of optimizing the task cost. Methods that optimize this cost produce trajectories resulting in end-effector pose errors much smaller than the ones obtained just sampling from the priors. E.g., in the EnvWarehouse-RobotPanda task, the end-effector mean error in orientation drops from $7.1$ to $1.2$ between Dprior and MPD.
@@
Figure 8: (Left) computation time and success rate, and (right) shortest path lengths results in the EnvSpheres3D-RobotPanda task (additional objects) across $100$ contexts. We compare one RRT Connect sample with cost optimization (RRTC + Cost), and MPD with increasing batches (11, $10$, $100$).
@@
Figure 1 presented a motivation for initializations in optimization-based planning methods. In our previous work [24], the experiments showed that (as expected) a sampling-based prior achieves $100\%$ success rates, but with more computation time (hence, we do not include it here in the main baselines). In this work, we consider sampling one trajectory using RRTConnect, since parallel sampling incurs too much computation time as it is not as easy to parallelize. We use the same implementation as described in section IV-A. The question we aim to answer is whether MPD has a benefit over first running a sampling-based planner and then solving the optimization problem from eq. 1 (we used the same number of steps as for MPD). For this experiment, we chose the EnvSpheres3D-RobotPanda task with additional objects because it is the hardest task we considered (due to narrow passages). Figure 8 shows the results obtained when planning $100$ contexts. We observe that sampling one trajectory with RRTConnect uses more computation time than sampling $100$ with a learned diffusion model. MPD’s success rate increases with the batch size, while computation time remains almost constant. The boxplot on the right side also shows that MPD produces trajectories with shorter lengths and less variance. This experiment shows that offloading sampling-based planner computations and compressing them into a trajectory diffusion model leads to faster and better planning results.
@@
i=Ni=NMPD denoising processi=0i=0
Figure 11: The figures illustrate the benefits of sampling with MPD instead of first sampling from the prior and then optimizing the cost function. LABEL:sub@fig:EnvEmpty2D_demonstrations The robot is a 22D point mass in an environment without any obstacles, and demonstration trajectories are obtained by sampling from a Gaussian Process prior. At inference, an obstacle (in red) is inserted into the environment. LABEL:sub@fig:EnvEmpty2D_diffusion_prior Samples from the learned diffusion prior. LABEL:sub@fig:EnvEmpty2D_diffusion_prior_then_guide Results of first sampling from the prior and then optimizing the collision cost function. LABEL:sub@fig:EnvEmpty2D_mpd Samples from the posterior distribution generated with MPD. The second row illustrates how MPD’s denoising process moves from Gaussian noise to clean trajectories while avoiding the red obstacle. Trajectories that are in collision with the red obstacles are shown in black, and collision-free trajectories in orange.
In this section, we answer Q3. One important aspect when using cost-guided diffusion in the context of motion planning is to understand in what situations can sampling from the denoising posterior distribution eq. 8 be beneficial, compared to first sampling from the diffusion prior and then optimizing the cost function starting from the proposed samples. In fig. 11 we show a simple example to motivate this difference. We consider a 22D point mass in an environment without any obstacle (fig. 11a), generate a small set of demonstrations with a Gaussian Process trajectory, and learn a trajectory diffusion prior. At inference we add a new obstacle, the red square, and do optimization using the object collision cost. The figures show two aspects of MPD. First, while sampling from the diffusion prior and then optimizing the cost fails to provide collision-free trajectories (fig. 11c), MPD can obtain a larger percentage of those (fig. 11d). This phenomenon happens because once a large portion of the trajectory is in collision with the obstacle, it becomes more difficult to remove it from collision. Second, in fig. 11d, we notice how trajectory samples from MPD evolve during the denoising process. The evolution shows that sampling from the posterior distribution moves the samples towards regions of the configuration space where the posterior has higher probability density, thus generating samples that are collision-free but close to the prior distribution. Therefore, the benefits of sampling from the posterior are more diverse and collision-free trajectories, which can be observed in fig. 6 by comparing Dprior+Cost and MPD results, especially in the environments with additional objects.
@@ IV-F Impact of the B-spline Parametrization
One of this work’s contributions is performing diffusion in the space of B-spline coefficients instead of waypoints. Hence, to answer Q4, we compare both representations. The benefits have been detailed in section III-E, but perhaps the main one is generating smooth trajectories. As the velocity and acceleration of a linear waypoint trajectory representation are computed with central finite differences, they are constant between waypoints and lead to large jumps. There are two ways to get smoother trajectories with this representation. First, we can increase the number of waypoints, as commonly done in optimization-based algorithms. Second, after running the denoising process, we can do extra optimization steps to smoothen and optimize the trajectory to remain collision-free and within joint limits while minimizing the task cost. Both of these options have drawbacks in terms of additional computations, e.g., in the first case, the denoising network needs to process larger inputs. We performed an ablation experiment where we replaced the B-spline with waypoints. We used the same number of waypoints as B-spline coefficients, which guarantees that the denoising network size and computations are similar. As expected, there are no major differences across all metrics except for smoothness values  [^1] Smoothness is measured as the sum of accelerations along the trajectory., which are better (lower) for the B-spline compared to the sparse waypoint representation. To better grasp the difference visually, fig. 12 shows the distribution of smoothness values obtained with MPD across all contexts and tasks. The waypoint representation consistently leads to worse smoothness. Figure 13 shows a visual comparison in the EnvSimple2D-RobotPointMass2D tasks.
Figure 14: Computation times for denoising a batch of $1000$ trajectories with B-splines and dense Waypoints using $H=128$ and $d=7$ with $15$ steps of DDIM.
The figures show that even though the waypoint representation can generate valid trajectories, they are not smooth, as can be seen in the straighter path in the environment plot and in the velocity and acceleration time profiles. These types of trajectories are harder for a robot controller to follow and can lead to jerky movements.
A benefit of B-splines in comparison to using a dense trajectory is the faster generation time, as the denoising function processes inputs with lower dimensions. To generate trajectories with $H$ points, we can sample from a model that outputs this trajectory size, or we can sample $n_{b}<H$ control points of a B-spline, and then interpolate the trajectory to $H$ using eq. 19. Figure 14 shows the computation times when generating a batch of trajectories with B-splines for an increasing number of control points, resulting from sampling the diffusion model and interpolating with eq. 19, and sampling waypoint trajectories with $H=128$. We consider only denoising times and interpolation (whose time is minimal), since the computation of costs/gradients only depends on the number of dense points $H$, and is the same for B-splines or dense waypoint representations. The results show the computational benefit of performing diffusion with lower-dimensional representations, instead of dense trajectories.
@@ IV-G Learning and Adapting from Human Demonstrations
Learning the prior distribution on trajectories is agnostic to the expert demonstrations. Previously, we assumed that trajectories would be generated by a sampling-based path planner. However, in real-world tasks, it is sometimes desirable and easier for a human to provide demonstrations via kinesthetic teaching of what movement they would like the robot to execute. Therefore, to answer Q5, in this experiment, we learn the trajectory distribution directly from human demonstrations, showing that MPD can be used in this scenario as well. Using the EnvWarehouse-RobotPanda environment, we consider the task of moving one item (a tea box) from several locations in the small to the big shelf and letting a user perform multiple demonstrations from random start and goal poses by moving the robot in gravity compensation with kinesthetic teaching from regions on the small shelf to the big shelf, generating approximately $100$ trajectories. These are augmented using small noise values to perform a total of approximately 44k trajectories. Examples of human-demonstrated trajectories are shown in fig. 15. The diffusion prior model is then trained on joint space trajectories similarly and with the same hyperparameters as the simulation experiments.